OpenStudy (anonymous):

for the following equation find the discriminant and describe nature of the solutions -14x^2-7x+17

OpenStudy (anonymous):

i know ;)

OpenStudy (anonymous):

the discriminat is the radicand from the quadratic formula\[b^2-4ac\]where\[ax^2+bx+c=0\] in you problem\[b^2-4ac=49-4(-14)(17)=1001\]which is positve but not a perfect square, so there are "two real roots"

OpenStudy (anonymous):

ok, i was sure if i had the radicand thing right, thanks!

OpenStudy (anonymous):

assuming a, b, c are rational numbers b^2-4ac Number & Nature of Roots ====== ================ 0 one double real root >0 two real roots <0 two complex roots

OpenStudy (anonymous):

ur welcome :})